Let the probability of getting head for a biased coin be \[\frac{1}{4}\]. It is tossed repeatedly until a head appears. Let \[\mathrm{N}\]
Solutions · 1
Official solution
We have the quadratic equation 64x2+5Nx+1=0. For it to have no real roots, the discriminant (b2−4ac) should be less than 0. Here, a=64, b=5N, and c=1.
This gives us :
(5N)2−4×64×1<0⇒25N2<256⇒N2<25256⇒N<25256=516
Since N must be an integer (as it represents the number of tosses), the possible values of N are 1, 2, or 3.
The probability of getting the first head on the n-th toss (given the probability of getting a head is 1/4) is given by the geometric distribution formula, (1−p)n−1×p.
So, the probability for our specific values of N is:
P(N=1)=(1−1/4)1−1×(1/4)=1/4P(N=2)=(1−1/4)2−1×(1/4)=3/4×1/4=3/16P(N=3)=(1−1/4)3−1×(1/4)=(3/4)2×1/4=9/64
Therefore, the total probability (p/q) is :
p/q=P(N=1)+P(N=2)+P(N=3)=1/4+3/16+9/64=16/64+12/64+9/64=37/64
So, p=37, q=64 and q−p=64−37=27.
Therefore, q−p is equal to 27.